# Don’t be lazy: CompleteP enables compute-efficient deep transformers

### v2

@@ 5 Re-examining compute-optimal N:L ratio
We now turn to the study of the optimal $N\!:\!L$ ratio in transformers, as its tuning can significantly impact compute efficiency. Kaplan et al. [2] popularized the practice of fixing $N:L\approx 100$ and scaling both model parameters and tokens [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72], but without adopting a parameterization that allows for infinite width and depth control. The consequences of this are exemplified in Section 4, where we have shown that deeper models suffer under SP/µP due to hyperparameter detuning and instability, leading $\alpha=1$ to outperform. This depth-disadvantage represents an important confounding variable in these studies on the optimal $N\!:\!L$. In other words, without an adequate depth parameterization, deeper models are not given their best chance to succeed. In this section, we conduct the first study of compute-optimal $N\!:\!L$ with simultaneous control of infinite depth and width and show that $\alpha=1$ enables a wider range of close-to-compute-optimal aspect ratios, where even $N:L\approx 10$ remain close to compute-optimal.
@@
We train models in the compute-optimal setting of 20 TPP, select compute-efficient batch sizes based on FLOPs [2, 81, 82, 83], and ensure well-tuned $\eta,\lambda,\sigma$ for $L_{\text{base}}=2$ [84] as in the previous section. See Appendix I for further details. This time, we vary $N,L$ to approximately maintain the total number of non embedding parameters $P_{\text{non-emb}}=12N^{2}L$, for values of $P_{\text{non-emb}}\in\{50M,300M,1.5B\}$. This setup matches the model sizes from Kaplan et al. [2], but while they use the same number of tokens for all runs, we strictly follow the compute-optimal prescription of Hoffmann et al. [3]. See Section I.3 for extensive details on model shapes and plots.
@@ Results
Figure 4: Optimal $N$:$L$ across model sizes. (a)-(c) For models of size $\{50\text{M},300\text{M},1.5\text{B}\}$ we see an increase in optimal $N$:$L$ as size increases but optimal $N$:$L$ is less than $\approx 60$ in this range of model sizes. (d) Scaling laws with FLOPs for the optimal aspect ratios in each parameterization. (e) FLOP savings of CompleteP compared to $\mu$P baseline, as a function of depth $L$. (f) CompleteP ($\alpha=1$) enables a larger range of $N$:$L$ that are within $1\%$ of optimal. See Table 4 and Section I.3 for details.
@@
We first fit the power law of the form $\hat{X}(F)=(F/a)^{-b}$ describing how validation cross entropy loss ($X$) scales with a power law in FLOPs ($F$) [1, 2, 62]. The scaling law $\hat{X}(F)$, reported in Figure 4d is fitted with the most compute-efficient $N\!:\!L$ from each parameter count and it defines the compute-efficient frontier in our setup. To fairly compare model loss obtained with different FLOP budgets, we then evaluate each $N\!:\!L$ configuration in terms of the distance to the fitted $\hat{X}_{\alpha=1}(F)$. This approach is similar to Figure 5 from Kaplan et al. [2]; in our case it provides a measure of the loss difference with respect to the reference scaling law of models trained with CompleteP. We report $N\!:\!L$ as a function of this measure in Figure 4a-c. Notice that the optimal aspect ratio is the same in µP and CompleteP ($\alpha=1$), while $\alpha=0.5$ prefers slightly wider models. In all cases, CompleteP ($\alpha=1$) gives lower loss values. In particular, we observe the following trend, (see also Figure 4f) :
@@
Finding 3: Compute-optimal $N\!:\!L$ trends larger with increasing scale, for all parameterizations.
Advantages of CompleteP at large depth. The gap between $\mu$P and CompleteP increases for deeper models. We quantify this gap in terms of FLOPs savings in Figure 4e, where we can see:
@@
Finding 4: The deeper the model, the more FLOP savings CompleteP ($\alpha=1$) has over µP.
@@
We attribute this gap to the fact that at higher depth µP is more detuned due to lack of HP transfer. In the 1.5B parameter models, CompleteP saves 11.8% of FLOPs for optimal $N\!:\!L$ and 34.4% FLOPs in the deepest models (lowest $N\!:\!L$). Furthermore, we make the following observation:
@@
Finding 5: As model scale increases, CompleteP ($\alpha=1$) enables deep-narrow models (small $N\!:\!L$) to remain close to compute-optimal.
@@
For $P_{\text{non-emb}}$=1.5B, $\alpha=1$ enables $N:L=11.8$ to remain within 1% of compute-optimal, compared to $N:L=38.7$ from µP. Hardware latency worsens with increasing depth, making shallow-wide models preferable for latency-sensitive applications. However, low-memory hardware can benefit from narrow-deep models by streaming one layer at a time into memory. Prominent examples of weight streaming exist in both training [85] and inference [86, 87] settings.
@@
Downstream Performance. In Table 2 we evaluate the $P_{\text{non-emb}}$=1.5B models at the minimum and optimal $N\!:\!L$ settings, and confirm the upstream gains also translate to gains across five downstream tasks: HellaSwag [88] ARC-Easy [89], LAMBADA [90], RACE [91], PIQA [92], and BoolQ [93]. These collectively test for common sense reasoning, world knowledge, and reading comprehension. The details of downstream evaluation setup are in Section I.4.
@@
Table 2: Zero-shot downstream evaluation accuracy for $P_{\text{non-emb}}\approx$1.5B models at the optimal and minimum $N\!:\!L$ ratios. We report average task accuracy $\pm$ standard error. If the average task accuracy is within the standard error, both columns are bolded. The final row reports average accuracy $\pm$ pooled standard error across all tasks.
@@ 6 Desiderata for Hyperparameter Transfer
to ensure that the change in $\mathbf{h}^{\ell}$ entries is $\Theta(1)$ for any depth or width (see Appendix D.2 or [8]). We focus our presentation on the AdamW optimizer because it is widely used for LLM training, but the desiderata below can be equally well applied to any optimizer (see [6, 7, 8]).
@@
This desideratum prescribes how to scale weight variances (or pre-factors) as in Table 1. It also constrains the value of $\alpha$ to be at least $0.5$ (see Section D.1). It can be viewed as a numerical stability condition and has been well-studied in the signal-propagation literature [11, 16, 17, 19, 20, 21, 40]. We review why this desideratum imposes $\alpha\geq 1/2$ in Section D.1.
@@
Many parameterization schemes, including those from works on signal propagation [16, 17, 27, 12], the NTK parameterization, and the Mean Field / $\mu$P approaches [29, 33, 4], satisfy Desideratum 1 for width scaling $N\to\infty$ at fixed $L$. The core distinction between them is the presence or absence of feature learning. While there are several ways to make this precise, we follow the original $\mu$P work [4] and formalize feature learning by considering the change in hidden layer representations after each step of training.
@@
To derive a parameterization for HP transfer across width, [4] required not only that the relative change $||\Delta\mathbf{h}^{\ell+1}||_{2}/||\mathbf{h}^{\ell+1}||_{2}$ to pre-activations in layer $\ell$ be $\Theta(1)$ but also that the proportion of this change $||\Delta\mathbf{W}^{\ell}\phi(\mathbf{h}^{\ell})||_{2}/||\mathbf{h}^{\ell+1}||_{2}$ attributable directly to the parameters $W_{\ell}$ must be $\Theta(1)$ as well. This excludes degenerate situations such as when one trains only the first few layers of a network.
While the maximal update requirement leads to a unique parameterization for HP transfer across width in a fixed-depth network, it is not directly applicable to the setting where one also scales the network depth. For example, in the case when $\alpha=0.5$ and $\mathcal{F}_{\ell}(\mathbf{h}^{\ell};\bm{\theta}^{\ell})=\mathbf{W}^{\ell}\phi(\mathbf{h}^{\ell})$, one must actually require $\|\Delta\mathbf{W}^{\ell}\phi(\mathbf{h}^{\ell})\|^{2}/||\mathbf{h}^{\ell+1}||_{2}^{2}=\Theta(L^{-0.5})$ in order for activations to remain stable in the sense that $||\Delta\mathbf{h}^{\ell+1}||/||\mathbf{h}^{\ell+1}||=\Theta(1)$ (see Section D.2 & [6, 7]). To cover residual networks of growing depth, we therefore require that the maximal update prescription hold per residual block.
@@
Desideratum 2 matches the maximal update prescription of [4] when depth $L$ is finite, as $\Theta(1/L)=\Theta(1)$. Furthermore, Desideratum 2 uniquely determines the depth-dependence of the learning rate in AdamW to be $\eta=\Theta(L^{1-\alpha})$ for the update to satisfy the $\Theta(1/L)$ scale [6]. Note also that requiring $\mathcal{F}(h^{\ell};\theta_{\ell})=\Theta(1)$ with respect to $L$ constrains $\alpha\leq 1$. When combined with Desideratum 1 this explains why we consider $0.5\leq\alpha\leq 1$. We also emphasize that additional care must be taken to correctly determine learning rates for biases and LayerNorm (LN) parameters, otherwise $\alpha=0.5$ fails Desideratum 2 for a pre-LN transformer (see Figure 7 and Appendix E).
@@
$\mathbf{h}^{\text{lin},\theta}(\bm{\theta},\bm{\theta}_{0})=\mathbf{h}(\bm{\theta}_{0})+\langle\nabla_{\theta}\mathbf{h}(\theta)|_{\bm{\theta}_{0}},\bm{\theta}-\bm{\theta}_{0}\rangle.$
@@
Figure 6: Only $\alpha=1$ achieves complete feature learning with stability. SP/$\mu$P diverges at large depth (triangles are NaNs), while $\alpha=0.5$ converges to the linearization at a rate of $1/\sqrt{L}$. The y axis is the left hand side of Eq. 4.
@@
Note the first term is exactly the change $\Delta h_{\ell+1}^{\text{lin},\theta_{\ell}}$ to the linearization of $h^{\ell+1}$, which has the correct order $\Theta(L^{-1})$. The second term therefore captures the difference between updating $\mathbf{h}^{\ell+1}$ and $\mathbf{h}_{\ell+1}^{\text{lin},\theta_{\ell}}$ and has order $\Theta(L^{\alpha-2})$. The two terms have the same order only when $\alpha=1$. This implies that for all $\alpha<1$, as we increase the depth $L$ of the network, the contribution to the training dynamics of $\mathbf{h}^{\ell+1}$ of the non-linear term diminishes with scale. We empirically verify this, for the same toy model, in Figure 6 (see Section I.5 for details). Since we expect the optimal HPs in shallow models to depend on linear and non-linear dynamics of $\mathbf{h}_{\ell+1}$, we do not expect these HPs to be optimal when scaling depth with $\alpha<1$.
While we carried out our computations above only in the simple toy model (Equation Equation 5), the form of the calculation extends to any arbitrary $\mathcal{F}_{\ell}(\mathbf{h}^{\ell};\bm{\theta}^{\ell})$ with bounded depth, including large wide MLPs and self-attention blocks. The core observation is that higher order terms in the Taylor expansion ($k>1$) of $\Delta_{\bm{\theta}^{\ell}}\mathbf{h}^{\ell+1}$ will always have the form
@@
Finding 6: Only CompleteP ($\alpha=1$) ensures stable training, maximal updates, and complete feature learning as $N\to\infty,L\to\infty$.
@@ 7 Conclusion
We studied in large pre-LN transformers a variety of parameterizations, i.e., prescriptions for adjusting hyperparameters such as learning rates as functions of network architecture. Among them, we found that CompleteP ($\alpha=1$) gives HP transfer when varying depth and width. We showed, moreover, that relative to other proposed parameterizations, CompleteP realizes significant FLOP savings during pre-training, even in previously unexplored, approximately compute-optimal settings where batch size and dataset size grow together with model size. To achieve this, we extended the theoretical analysis of CompleteP to include prescriptions for how to scale LN parameters and AdamW’s $\epsilon$ across depth and width.
@@ Appendix A Broader impacts
Appendix B Limitations
This work has both empirical and theoretical limitations. Our empirical results only test the setting of next-token prediction decoder-only pre-LN transformers trained with AdamW on text data. This may limit the generality of our empirical results to other settings, but we believe our experimental setup is representative of other settings as well. We were also limited in the scale we could test due to our compute constraints. For example the $\hat{X}$ fits in Section 5 were only fit to 3 points due to compute constraints with training models larger than 1.5B parameters. These empirical $\hat{X}$ fits are known to be sensitive to small scale variability: for example if the smallest data point is less FLOP efficient due to having a larger proportion of embedding FLOPs, the fit can have a much steeper slope than is true at larger scales.
Our theoretical analysis considers scaling width and depth at fixed token count, which limits the direct applicability to the fixed TPP compute-optimal regime. However, the scaling predictions derived in the fixed token setting have predicted success in the fixed TPP compute-optimal regime.
Appendix C Coordinate check test for verification of stable training
To verify our theoretical expectations for training stability match empirical results, we perform “coordinate check” tests where we scale $L$ and measure the change in activation size at the final residual addition $h_{L}$ in Figure 7. Interestingly, $\alpha=0.5$ as described in the literature [6, 7, 8] did not achieve transformer training stability as we initially expected (2nd column). We reasoned this was due to a lack of consideration for how the bias and LayerNorm parameters can affect training stability. After adopting the bias and LayerNorm $\eta$ adjustments proposed in Table 1 and Appendix E, we achieved training stability with $\alpha=0.5$ (3rd column). Figure 7 shows both $\alpha\in\{0.5,1\}$ achieve stability for any depth. This test is a necessary but not sufficient condition for satisfying Desiderata 2.
Figure 7: Coordinate check. Frobenius norm of activations after merged residual streams from MHSA and MLP blocks across models of increasing depths after training step $t$. 1st column: SP/$\mu$P without depth control. 2nd column: $\alpha=0.5$ as described in literature [6, 7, 8]. 3rd column: $\alpha=0.5$ (Table 1). 4th column: $\alpha=1.0$. $\alpha\in\{0.5,1\}$ achieve stability for any depth.
Appendix D Scalings that Guarantee Our Desiderata

### v3

@@ 5 Re-examining compute-optimal N:L ratio
We now turn to the study of the optimal N:LN\!:\!L ratio in transformers, as its tuning can significantly impact compute efficiency. Kaplan et al. [2] popularized the practice of fixing $N:L\approx 100$ and scaling both model parameters and tokens [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72], but without adopting a parameterization that allows for infinite width and depth control. The consequences of this are exemplified in Section˜4, where we have shown that deeper models suffer under SP/µP due to hyperparameter detuning and instability, leading $\alpha=1$ to outperform. This depth-disadvantage represents an important confounding variable in these studies on the optimal N:LN\!:\!L. In other words, without an adequate depth parameterization, deeper models are not given their best chance to succeed. In this section, we conduct the first study of compute-optimal N:LN\!:\!L with simultaneous control of infinite depth and width and show that $\alpha=1$ enables a wider range of close-to-compute-optimal aspect ratios, where even $N:L\approx 10$ remain close to compute-optimal.
@@
We train models in the compute-optimal setting of 20 TPP, select compute-efficient batch sizes based on FLOPs [2, 86, 87, 88], and ensure well-tuned $\eta,\lambda,\sigma$ for $L_{\text{base}}=2$ [89] as in the previous section. See Appendix˜G for extensive experimental details and plotting methodology. This time, we vary $N,L$ to approximately maintain the total number of non embedding parameters $P_{\text{non-emb}}=12N^{2}L$, for values of $P_{\text{non-emb}}\in\{50M,300M,1.5B\}$. This setup matches the model sizes from Kaplan et al. [2], but while they use the same number of tokens for all runs (compute-inefficient), we follow the compute-optimal prescription of Hoffmann et al. [3]. Instead of pure web text [2], we train on SlimPajama: a mix of web text, academic prose, and code. We use the maximal update parameterization meaning our HPs have superior tuning with respect to changing width compared to Kaplan et al. [2] who use the standard parameterization. These differences modernize our setup and make our results more applicable to contemporary LLM training.
@@ Results
[…]
@@
We first fit the power law of the form $\hat{X}(F)=(F/a)^{-b}$ describing how validation cross entropy loss ($X$) scales with a power law in FLOPs ($F$) [1, 2, 62]. The scaling law $\hat{X}(F)$, reported in Figure˜4d is fitted with the most compute-efficient N:LN\!:\!L from each parameter count and it defines the compute-efficient frontier in our setup. To fairly compare model loss obtained with different FLOP budgets, we then evaluate each N:LN\!:\!L configuration in terms of the distance to the fitted $\hat{X}_{\alpha=1}(F)$. This approach is similar to Figure 5 from Kaplan et al. [2]; in our case it provides a measure of the loss difference with respect to the reference scaling law of models trained with CompleteP. We report N:LN\!:\!L as a function of this measure in Figure˜4a-c. Notice that the optimal aspect ratio is the same in µP and CompleteP ($\alpha=1$), while $\alpha=0.5$ prefers slightly wider models. In all cases, CompleteP ($\alpha=1$) gives lower loss values. In particular, we observe the following trend, (see also Figure˜4f) :
@@
Finding 4: Compute-optimal N:LN\!:\!L trends larger with increasing scale, for all parameterizations.
[…]
@@
Finding 5: The deeper the model, the more FLOP savings CompleteP ($\alpha=1$) has over µP.
@@
We attribute this gap to the fact that at higher depth µP is more detuned due to lack of HP transfer. In the 1.5B parameter models, CompleteP saves 11.8% of FLOPs for optimal N:LN\!:\!L and 34.4% FLOPs in the deepest models (lowest N:LN\!:\!L). The shaded regions in Figure 1-right and 4f represent the N:L range with $\leq 1\%$ loss increase relative to compute-optimal N:L (see Section˜G.2 for more detail).
@@
Finding 6: As model scale increases, CompleteP ($\alpha=1$) enables deep-narrow models (small N:LN\!:\!L) to remain close to compute-optimal.
@@
For $P_{\text{non-emb}}$=1.5B, $\alpha=1$ enables NN:$L=11.8$ to remain within 1% of compute-optimal, compared to NN:$L=38.7$ from µP. Hardware latency worsens with increasing depth, making shallow-wide models preferable for latency-sensitive applications. However, low-memory hardware can benefit from narrow-deep models by streaming one layer at a time into memory. Prominent examples of weight streaming exist in both training [90] and inference [91, 92] settings.
@@
Downstream Performance. In Table˜2 we evaluate the 20 TPP $P_{\text{non-emb}}$=1.5B models at the minimum and optimal N:LN\!:\!L settings, and confirm the upstream gains also translate to gains across five downstream tasks [93, 94, 95, 96, 97, 98] that collectively test for common sense reasoning, world knowledge, and reading comprehension. The details of downstream evaluation setup are in Section˜G.3.
@@
Table 2: Zero-shot downstream evaluation accuracy for 20 TPP $P_{\text{non-emb}}\approx$1.5B models at the optimal and minimum N:LN\!:\!L ratios. We report average task accuracy $\pm$ standard error. If the average task accuracy is within the standard error, both columns are bolded.
@@
Given the popularity of TPP>20 training for enhanced parameter efficiency [56], we also compare compare parameterizations at 200 TPP for $P_{\text{non-emb}}\in\{50M,300M\}$ in Table˜3. CompleteP ($\alpha=1$) consistently achieves the best loss across all parameterizations, even in the 200 TPP regime.
Table 3: SlimPJ validation loss for $P_{\text{non-emb}}\in\{50M,300M\}$ models trained for 200 TPP.
2.867 2.874 2.859 2.973 2.998 2.950 2.451 2.455 2.443 2.537 2.534 2.497
@@ 6 Desiderata for Hyperparameter Transfer
to ensure that the change in $\mathbf{h}^{\ell}$ entries is $\Theta(1)$ for any depth or width (see Appendix C.2 or [8]). We focus our presentation on the AdamW optimizer because it is widely used for LLM training, but the desiderata below can be equally well applied to any optimizer (see [6, 7, 8]).
@@
This desideratum prescribes how to scale weight variances (or pre-factors) as in Table˜1. It also constrains the value of $\alpha$ to be at least $0.5$ (see Section˜C.1). It can be viewed as a numerical stability condition and has been well-studied in the signal-propagation literature [11, 16, 17, 19, 20, 21, 40]. We review why this desideratum imposes $\alpha\geq 1/2$ in Section˜C.1.
@@
[…]
@@
[…]
@@
Desideratum 2 matches the maximal update prescription of [4] when depth $L$ is finite, as $\Theta(1/L)=\Theta(1)$. Furthermore, Desideratum 2 uniquely determines the depth-dependence of the learning rate in AdamW to be $\eta=\Theta(L^{1-\alpha})$ for the update to satisfy the $\Theta(1/L)$ scale [6]. Note also that requiring $\mathcal{F}(h^{\ell};\theta_{\ell})=\Theta(1)$ with respect to $L$ constrains $\alpha\leq 1$. When combined with Desideratum 1 this explains why we consider $0.5\leq\alpha\leq 1$. We also emphasize that additional care must be taken to correctly determine learning rates for biases and LayerNorm (LN) parameters, otherwise $\alpha=0.5$ fails Desideratum 2 for a pre-LN transformer (see Figure˜7 and Appendix˜D).
@@
$\mathbf{h}^{\text{lin},\theta}(\bm{\theta},\bm{\theta}_{0})=\mathbf{h}(\bm{\theta}_{0})+\langle\nabla_{\bm{\theta}}\mathbf{h}(\bm{\theta})|_{\bm{\theta}_{0}},\bm{\theta}-\bm{\theta}_{0}\rangle.$
@@
[…]
@@
Note the first term is exactly the change $\Delta h_{\ell+1}^{\text{lin},\theta_{\ell}}$ to the linearization of $h^{\ell+1}$, which has the correct order $\Theta(L^{-1})$. The second term therefore captures the difference between updating $\mathbf{h}^{\ell+1}$ and $\mathbf{h}_{\ell+1}^{\text{lin},\theta_{\ell}}$ and has order $\Theta(L^{\alpha-2})$. The two terms have the same order only when $\alpha=1$. For all $\alpha<1$, as we increase the depth $L$ of the network, the contribution of $\mathbf{h}^{\ell+1}$ to the non-linear term diminishes with scale. We empirically verify this, for the same toy model, in Figure˜6 (see Section˜G.4 for details). Since we expect the optimal HPs in shallow models to depend on linear and non-linear dynamics of $\mathbf{h}_{\ell+1}$, we do not expect these HPs to be optimal when scaling depth with $\alpha<1$.
While we carried out our computations above only in the simple toy model (Equation Equation˜5), the form of the calculation extends to any arbitrary $\mathcal{F}_{\ell}(\mathbf{h}^{\ell};\bm{\theta}^{\ell})$ with bounded depth, including large wide MLPs and self-attention blocks. The core observation is that higher order terms in the Taylor expansion ($k>1$) of $\Delta_{\bm{\theta}^{\ell}}\mathbf{h}^{\ell+1}$ will always have the form
@@
Finding 7: Only CompleteP ($\alpha=1$) ensures stable training, maximal updates, and complete feature learning as $N\to\infty,L\to\infty$.
@@
7 Limitations
Our theoretical analysis considers scaling width and depth at fixed token count, which limits the direct applicability to the fixed TPP compute-optimal regime. However, the scaling predictions derived in the fixed token setting (Figure˜2) have predicted success in the fixed TPP compute-optimal regime (Figure˜3). The $\hat{X}$ fits in Section˜5 were only fit to 3 points due to budget constraints with training models larger than 1.5B parameters. However the three points span a 256x FLOP range and are each the most compute-efficient point from each group of 7-10 N:L values for each parameter count (Table˜5), so they are likely to be an accurate picture of the compute-efficient frontier. We use the scaling law fits for interpolation rather than extrapolation, where fitting to limited data can make predictions unreliable.
@@ 8 Conclusion
We studied in large pre-LN transformers a variety of parameterizations, i.e., prescriptions for adjusting hyperparameters such as learning rates as functions of network architecture. Among them, we found that CompleteP ($\alpha=1$) gives HP transfer when varying depth and width. We showed that CompleteP achieves significant FLOP savings during pre-training compared to other parameterizations, even in compute-optimal settings with jointly scaled batch, dataset, and model sizes. To achieve this, we extended the theoretical analysis of CompleteP to include prescriptions for how to scale LN parameters and AdamW’s $\epsilon$ across depth and width.
@@ Appendix A Broader impacts
Appendix B Coordinate check test for verification of stable training
To verify our theoretical expectations for training stability match empirical results, we perform “coordinate check” tests where we scale $L$ and measure the change in activation size at the final residual addition $h_{L}$ in Figure˜7. Interestingly, $\alpha=0.5$ as described in the literature [6, 7, 8] did not achieve transformer training stability as we initially expected (2nd column). We reasoned this was due to a lack of consideration for how the bias and LayerNorm parameters can affect training stability. After adopting the bias and LayerNorm $\eta$ adjustments proposed in Table˜1 and Appendix˜D, we achieved training stability with $\alpha=0.5$ (3rd column). Figure˜7 shows both $\alpha\in\{0.5,1\}$ achieve stability for any depth. This test is a necessary but not sufficient condition for satisfying Desiderata 2.
[…]
Appendix C Scalings that Guarantee Our Desiderata